OpenStudy (anonymous):

solve the inequality:

OpenStudy (anonymous):

\[\left| 7-2x \right|\ge3\]

OpenStudy (anonymous):

i got 5 ___ x ____ 2 i just dont know what signs to put

OpenStudy (anonymous):

\[|2x-7|\geq 3\] \[2x-7\leq -3\] or \[2x-7\geq 3\] solve each separately

OpenStudy (anonymous):

\[2x-7\leq -3\] \[2x\leq 4\] \[x\leq 2\] or \[2x-7\geq 3\] \[2x\geq 10\] \[x\geq 5\]